Sulfuryl chloride decomposes at high temperatures to produce sulfur dioxide and chlorine gases:SO2Cl2(g) SO2(g) + Cl2(g)At 375C, the equilibrium constant Kcis 0.045. If there are 2.0 grams of sulfuryl chloride, 0.17 gram of sulfur dioxide, and 0.19 gram of chlorine present in a 1.0 Liter flask,a) What is the value of the reaction quotient?b) Is the system at equilibrium?c) In which direction will the system move to reach equilibrium?
Answer #1

moles of SO2Cl2 = 2.0 / 134.97 = 0.0148 mol

moles of SO2 = 0.17 / 64.06 = 2.65 x 10^-3 mol

moles of Cl2 = 0.19 / 70.906 = 2.68 x 10^-3 mol

So2Cl2 (g)   <----------> SO2 (g) + Cl2 (g)

a)

Qc = [SO2] [Cl2] / [SO2Cl2]

    = (2.65 x 10^-3)( 2.68 x 10^-3) / (0.0148)

Qc = 4.8 x 10^-4 ,

reaction quotient = 4.8 x 10^-4

b)

Kc is not equal to Qc . so

no this system is not at equilibrium

c)

Kc > Qc

system move to right direction (forward direction)